OpenStudy (anonymous):

Write an equation for the line w/ points (-5,-1) (2,2) and has a slope of 3/7

OpenStudy (anonymous):

y=(3/7)x+b 2=(3/7)2+b 2=(6/7)+b (14/7)-(6/7)=(6/7)-(6/7)+b (8/7)=b y=(3/7)x+(8/7)